OpenStudy (anonymous):

How many roots does 2x^2+5x+2 have?

OpenStudy (anonymous):

Two postive?

OpenStudy (blurbendy):

(2x^2 + 4x + 2) = (x + 2)(2x + 1) x = -2 x = -1/2 2 roots

OpenStudy (anonymous):

Are they postive right?

OpenStudy (blurbendy):

the roots are negative

OpenStudy (anonymous):

Both Negative -2.-1/2 right?

OpenStudy (blurbendy):

yes

OpenStudy (anonymous):

Thanks :D

OpenStudy (blurbendy):

np